# MoC-System: Efficient Fault Tolerance for Sparse Mixture-of-Experts Model Training

### v2

@@ 3. Partial Experts Checkpointing
Refer to caption (a) PLT (%) Refer to caption (b) Validation loss
@@ 3.1.2. Impact on Model Accuracy
where $N_{moe}$ denotes the number of MoE layers within the model, and $N_{fault}$ denotes the count of faults encountered during the training. $L_{i,j}$ refers to the measured number of the $ith$ MoE layer’s lost tokens caused by the $jth$ fault, which is influenced by the checkpointing interval $I_{ckpt}$, the number $K_{pec}$ of saved experts, and the function $F$ for partial experts selection (e.g. sequential or load-aware methods). The product of the number of input tokens $T_{i}$ and $TopK_{i}$ of MoE gating indicates the total number of tokens processed by all experts in the $ith$ MoE layer during the training. It is worth noting that the actual count of tokens processed by all experts is typically less than $T_{i}\cdot TopK_{i}$, primarily due to the token dropout imposed by the expert capacity (lepikhin2020gshard, ).
To investigate the correlation between PLT and model accuracy, we conduct experiments that train GPT-125M-8E models with varying PEC configurations (different values of $K_{pec}$ and $I_{ckpt}$) on Wikitext dataset (wikitext, ). Each model’s training process is designed to encounter a fault at the midpoint, followed by a recovery from the saved PEC checkpoint.
@@ 3.2. Partial Experts Selection
More importantly, considering that experts within each MoE layer are distributed across various ranks and devices via EP, as depicted in Figure 4, the partial experts selection significantly affects the workload distribution across ranks, thus impacting the time cost of checkpointing. The most imbalanced workload scenario, for instance, involves checkpointing “Expert(1-0, 3-0, 5-0, 7-0)”, all located in “Rank0”. In this case, the checkpointing duration is primarily prolonged by “Rank0”, which bears the heaviest workload, while the other ranks remain underutilized.
@@
Load-aware Selection. We extend our investigation into the function of partial experts selection by incorporating a load-aware approach that prioritizes the checkpointing of $K_{pec}$ experts, characterized by the highest number of unsaved updates. Based on our empirical results, load-aware selection achieves model accuracy on par with sequential selection but necessitates more complicated control mechanisms and incurs higher costs. Consequently, we believe that sequential selection is a more practical method.
@@ 4. Fully Sharded Checkpointing
Figure 7. An illustration of two distinct checkpointing methods employed for training the MoE model, configured with DP = 4 and EP = 2. (a) illustrates the baseline method provided by the Megatron-DeepSpeed framework. (b) presents our proposed fully sharded checkpointing with equal sharding. For simplification, the model states are divided into two segments: the non-expert and the expert parts. The horizontal segments within each part represents various layers.
As discussed in Section 2.3.2, existing work lacks an efficient data-parallel sharding strategy for checkpointing MoE models in distributed training. Figure 7 (a) demonstrates that the baseline method provided by the Megatron-DeepSpeed framework (Megatron_DeepSpeed, ) only utilizes “Rank0” to save non-expert states and “EP-Group-0” to save expert states.
4.1. Equal Sharding for Expert Part
In contrast to the hybrid strategy depicted in Figure 1, which employs a single EP group, the prevailing practice (liu2024deepseek, ; wei2024skywork, ) for large-scale distributed training of MoE models employs multiple EP groups, as demonstrated in Figure 6.
@@ 4.2. Equal Sharding for Non-Expert Part
Given the considerable overhead associated with fine-grained sharding methods (nicolae2020deepfreeze, ) and the fact that the model parameters of the non-expert part comprise only 2% of the total checkpoint volume, we implement a coarse-grained sharding approach, utilizing layers (e.g. Attention and FFN) as the minimum partition units. Building upon this framework, we introduce an equal sharding strategy, aiming to evenly distribute the workload of checkpointing non-expert layers across all DP ranks, as depicted in Figure 7 (b). We define the ideal checkpointing workload of each rank, $C_{rank}$, using the following formulation:
@@ 4.3. Adaptive Sharding for Non-Expert Part
In our implementation, sharding strategies are exclusively utilized to partition model parameters, tailored to our specific scenario of ZeRO-2 DP + EP, where optimizer states are already partitioned, as depicted in Figure 6. Nevertheless, our methodologies are applicable to the partitioning of both model parameters and optimizer states in scenarios that do not incorporate ZeRO sharding (rajbhandari2020zero, ).
@@ 5.1. Two-level PEC Saving and Recovery
5.2. Asynchronous Checkpointing & Triple-Buffer
To minimize the overhead of states saving $O_{save}$, we implement an asynchronous checkpointing mechanism. This approach allows checkpointing to be overlapped with the normal training processes through the use of triple-buffer management. As illustrated in Figure 9, the triple-buffer comprises snapshot, persist, and recovery buffers, each meticulously designed to ensure data integrity during the saving process and data consistency during recovery. Initially, all three buffers are in the snapshot status. Each snapshot process involves the transfer of data from the GPU into one of these snapshot buffers. Upon the completion of a snapshot process—and in the absence of an ongoing persist buffer—the corresponding buffer transitions to the persist status, starting the transfer of data to persistent storage. Following the completion of the persist process, the buffer then becomes a recovery buffer, reflecting the latest checkpoint available for recovery, until another persist buffer transitions.
@@ 5.3. Adaptive Configuration for Two-Level PEC
Existing studies minimize $O_{lost}$ by reducing the checkpointing interval $I_{ckpt}$ (mohan2021checkfreq, ; wang2023gemini, ; eisenman2022check, ). This approach may increase $O_{save}\frac{I_{total}}{I_{ckpt}}$, necessitating a considered trade-off between the two metrics. In addition to $I_{ckpt}$, we introduce two new adjustable hyperparameters, $K_{snapshot}$ and $K_{persist}$, aimed at reducing the durations of snapshots and persistence, respectively. However, this advancement presents a new trade-off between efficiency and PLT.
To navigate these trade-offs across various software and hardware scenarios, we propose an adaptive configuration scheme for two-level PEC. Our primary strategy involves optimizing the value of $K_{snapshot}$ for snapshot-PEC to ensure it can be completely overlapped by the next F&B, thereby minimizing $O_{save}$ while achieving a low PLT. Even though the persist process can be fully overlapped with the subsequent training, its duration determines the lower bound for $I_{ckpt}$. As our tow-level recovery method significantly reduces the PLT caused by persist-PEC, $K_{persist}$ can be set to a relatively small value, which in turn, achieves the lowest $I_{ckpt}$.
@@
6. Evaluation
6.1. Experimental Setup
We implement our proposed MoC-System and conduct extensive experiments upon the Megatron-DeepSpeed (Megatron_DeepSpeed, ; smith2022using, ; rajbhandari2022deepspeed, ), which is an acclaimed open-source framework supporting the distributed training of MoE models. As shown in Table 1, we experiment with both language and vision models. The experimental language models (GPT-125M-8E and GPT-350M-16E) are extensions of GPT-3 like NLG model (brown2020language, ), provided by DeepSpeed-MoE (rajbhandari2022deepspeed, ). The GPT-125M-8E model is pre-trained on the Wikitext-2 dataset (wikitext, ) for the correlation analysis between PLT and final trained accuracy, as illustrated in Figure 5. The GPT-350M-16E model is pre-trained on a 1B token subset of the SlimPajama-627B dataset (cerebras2023slimpajama, ). Moreover, the distributed training of the GPT-350M-16E model is experimented with the three different configurations, as shown in Table 2. Using the hybrid parallel strategy of ZeRO-2 DP (rajbhandari2020zero, ) and EP, the training is deployed on a cluster that comprises a total of 60 nodes with 8×A800-SXM4-80GB GPUs each. The vision model, SwinV2-MoE (hwang2023tutel, ; swin, ), is trained on the ImageNet-1K dataset (deng2009imagenet, ).
(a) Total Checkpoint Size
(b) Bottleneck Rank in Case1
(c) Bottleneck Rank in Case2
(d) Bottleneck Rank in Case3
Figure 10. Experimental results of checkpoint size. (a) shows the impact of PEC on total checkpoint size. (b-d) illustrate the checkpointing workload of the bottleneck rank across various distributed configurations. ”EE” indicates equal sharding for the expert part, while ”EN” and ”AN” represent equal sharding and adaptive sharding for the non-expert part, respectively.
(a) Case1 Configuration
(b) Case2 Configuration
(c) Case3 Configuration
@@ 6.2.1. Checkpoint Size
We evaluate the effectiveness of PEC in reducing checkpoint size through experiments on the GPT-350M-16E model training. Unless otherwise specified in subsequent experiments, PEC employs the sequential selection strategy, and the baseline refers to the method provided by the Megatron-DeepSpeed framework (Megatron_DeepSpeed, ). As illustrated in Figure 10 (a), the total checkpoint size for each process decreases as $K_{pec}$ decreases, reaching 42.3% of the full model checkpoint size when $K_{pec}$ is set to 1.
However, merely reducing the total checkpoint size is insufficient for optimizing efficiency in distributed training. As the checkpointing workload is distributed across various training ranks, the duration of the blocking checkpointing process is primarily determined by the bottleneck rank, which has the heaviest workload and longest processing time. Therefore, we further assess the checkpointing workload of the bottleneck rank using different sharding strategies, as illustrated in Figure 10 (b-d).
The results indicate that our fully sharded checkpointing strategy, which applies equal sharding to both non-expert and expert parts, significantly reduce the workload of bottleneck rank in both full saving (12% to 28%) and PEC scenarios (22% to 29%). Notably, equal sharding of the expert part is only effective in scenarios with multiple EP groups (Case 3). With $K_{pec}=1$, adaptive sharding of the non-expert part can further reduce the workload by 3.7% to 6.1% compared to the equal sharding approach.
@@ 6.2.3. Asynchronous Checkpointing
Given that our approaches have been validated to reduce the duration of the checkpointing process, we further assess the end-to-end optimization efficacy of our MoC-System, which implements an asynchronous checkpointing process. As illustrated in Figure 12, the fully optimized asynchronous process in our MoC-System (MoC-Async) can decrease checkpointing overhead by 98.2% to 98.9% and accelerate each training iteration by 3.25 to 5.12 times in the three experimental cases, compared to the baseline using blocking checkpointing.
@@
Applying our asynchronous process directly to the baseline checkpointing method (Base-Async) can overlap 97.9% of the checkpointing time in Case 2, as its snapshot duration is sufficiently short to be fully overlapped. However, this method can only overlap 86.3% and 92.1% in Case 1 and Case 3, respectively, because their durations are too long to be fully overlapped. Employing all optimizations within the MoC-System can further achieve a 1.4% to 33.2% improvement over Base-Async in the three cases.
@@
In addition to reducing $O_{save}$, our MoC-Async achieves half the $I_{ckpt}$ compared to the Base-Async method, as it takes only half the time to complete the snapshot and persist process. For instance, it reduces $I_{ckpt}$ from 2.3 to 1.2 in Case 2. Consequently, our MoC-Async can minimize the overall checkpoint overhead $O_{ckpt}$.

### v3

@@ 3.1.2. Impact on Model Accuracy
where $N_{moe}$ denotes the number of MoE layers within the model, and $N_{fault}$ denotes the count of faults encountered during the training. $L_{i,j}$ refers to the measured number of the $ith$ MoE layer’s lost tokens caused by the $jth$ fault, which is influenced by the checkpointing interval $I_{ckpt}$, the number $K_{pec}$ of saved experts, and the function $F$ for partial experts selection (e.g. sequential or load-aware methods). The product of the number of input tokens $T_{i}$ and $TopK_{i}$ of MoE gating indicates the total number of tokens processed by all experts in the $ith$ MoE layer during the training. It is worth noting that the actual count of tokens processed by all experts is typically less than $T_{i}\cdot TopK_{i}$, primarily due to the token dropout imposed by the expert capacity (Lepikhin et al., 2020).
To investigate the correlation between PLT and model accuracy, we conduct experiments that train GPT-125M-8E models with varying PEC configurations (different values of $K_{pec}$ and $I_{ckpt}$) on Wikitext dataset (Merity et al., 2016). Each model’s training process is designed to encounter a fault at the midpoint, followed by a recovery from the saved PEC checkpoint.
@@ 3.2. Partial Experts Selection
More importantly, considering that experts within each MoE layer are distributed across various ranks and devices via EP, as depicted in Figure 5, the partial expert selection significantly affects the workload distribution across ranks, thus impacting the time cost of checkpointing. The most imbalanced workload scenario, for instance, involves checkpointing “Expert(1-0, 3-0, 5-0, 7-0)”, all located in “Rank0”. In this case, the checkpointing duration is primarily prolonged by “Rank0”, which bears the heaviest workload.
@@
Load-aware Selection. We extend our investigation into the function of partial expert selection by incorporating a load-aware approach that prioritizes the checkpointing of $K_{pec}$ experts, characterized by the highest number of unsaved updates. Based on our empirical results, load-aware selection achieves model accuracy on par with sequential selection but necessitates more complicated control mechanisms and incurs higher costs, making it a less favorable option.
@@ 4. Fully Sharded Checkpointing
As discussed in Section 2.3.2, existing work lacks an efficient data-parallel sharding strategy for checkpointing MoE models in distributed training. Figure LABEL:fig:fully_sharded_checkpointing_a demonstrates that the baseline method provided by the Megatron-DeepSpeed framework (Microsoft, 2022) only utilizes “Rank0” to save non-expert states and “EP-Group-0” to save expert states.
4.1. Equal Sharding for Expert Part
In contrast to the hybrid strategy depicted in Figure 2, which employs a single EP group, the prevailing practice (Liu et al., 2024a; Wei et al., 2024) for large-scale distributed training of MoE models employs multiple EP groups, as demonstrated in Figure 9.
@@ 4.2. Equal Sharding for Non-Expert Part
Given the considerable overhead associated with fine-grained sharding methods (Nicolae et al., 2020) and the fact that the model parameters of the non-expert part comprise only 2% of the total checkpoint volume, we implement a coarse-grained sharding approach, utilizing layers (e.g. Attention and FFN) as the minimum partition units. Building upon this framework, we introduce an equal sharding strategy, aiming to evenly distribute the workload of checkpointing non-expert layers across all DP ranks, as depicted in Figure LABEL:fig:fully_sharded_checkpointing_b. We define the ideal checkpointing workload of each rank, $C_{rank}$, using the following formulation:
@@ 4.3. Adaptive Sharding for Non-Expert Part
In our implementation, sharding strategies are exclusively utilized to partition model parameters, tailored to our specific scenario of ZeRO-2 DP + EP, where optimizer states are already partitioned, as depicted in Figure 9. Nevertheless, our methodologies are applicable to the partitioning of both model parameters and optimizer states in scenarios that do not incorporate ZeRO sharding (Rajbhandari et al., 2020).
Figure 11. An illustration of two distinct checkpointing methods employed for training the MoE model, configured with DP = 4 and EP = 2. (a) illustrates the baseline method provided by the Megatron-DeepSpeed framework. (b) presents our proposed fully sharded checkpointing with equal sharding. For simplification, the model states are divided into two segments: the non-expert and the expert parts. The horizontal segments within each part represent various layers.
@@ 5.1. Two-level PEC Saving and Recovery
5.2. Asynchronous Checkpointing & Triple Buffering
To minimize the overhead of states saving $O_{save}$, we implement an asynchronous checkpointing mechanism that allows checkpointing to overlap with the normal training processes. Specifically, we develop an agent at each node to facilitate the two-level checkpointing management through a triple-buffer mechanism. As illustrated in Figure 13, the triple buffering comprises snapshot, persist, and recovery buffers, each meticulously designed to ensure data integrity during the saving process and data consistency during recovery. Initially, all of these buffers are in the snapshot status. Each snapshot process, initiated by the asynchronous thread within each training process, involves the transfer and serialization of data from the GPU into one of these snapshot buffers. Upon the completion of a snapshot process—and in the absence of an ongoing persist buffer—the corresponding buffer transitions to the persist status, starting the transfer of data to persistent storage. Following the completion of the persist process, the buffer then becomes a recovery buffer, reflecting the latest checkpoint available for recovery, until another persist buffer transitions.
@@ 5.3. Adaptive Configuration for Two-Level PEC
Existing studies minimize $O_{lost}$ by reducing the checkpointing interval $I_{ckpt}$ (Mohan et al., 2021; Wang et al., 2023a; Eisenman et al., 2022). This approach may increase $O_{save}\frac{I_{total}}{I_{ckpt}}$, necessitating a considered trade-off between the two metrics. In addition to $I_{ckpt}$, we introduce two new adjustable hyperparameters, $K_{snapshot}$ and $K_{persist}$, aimed at reducing the durations of snapshot and persist, respectively, presenting a new trade-off between efficiency and PLT.
To navigate these trade-offs across various software and hardware scenarios, we propose an adaptive configuration scheme for two-level PEC. Our primary strategy involves optimizing the value of $K_{snapshot}$ for snapshot-PEC to ensure it can be completely overlapped by the next F&B, thereby minimizing $O_{save}$ while achieving a low PLT. Even though the persist process can be fully overlapped with the subsequent training, its duration determines the lower bound for $I_{ckpt}$. As our two-level recovery method significantly reduces the PLT caused by persist-PEC, $K_{persist}$ can be set to a relatively small value, which in turn, achieves the lowest $I_{ckpt}$.
Figure 17. Experimental results of checkpoint size. (a) shows the impact of PEC on total checkpoint size. (b-d) illustrate the checkpointing workload of the bottleneck rank across various distributed configurations. “EE” indicates equal sharding for the expert part, while “EN” and “AN” represent equal sharding and adaptive sharding for the non-expert part, respectively.
@@
6. Evaluation
6.1. Experimental Setup
We implement our proposed MoC-System and conduct extensive experiments upon the Megatron-DeepSpeed (Microsoft, 2022; Smith et al., 2022; Rajbhandari et al., 2022), which is an acclaimed open-source framework supporting the distributed training of MoE models. As shown in Table 1, we experiment with both language and vision models. The experimental language models (GPT-125M-8E and GPT-350M-16E) are extensions of GPT-3 like NLG model (Brown et al., 2020), provided by DeepSpeed-MoE (Rajbhandari et al., 2022). The GPT-125M-8E model is pre-trained on the Wikitext-2 dataset (Merity et al., 2016) for the correlation analysis between PLT and final trained accuracy, as illustrated in Figure 7. The GPT-350M-16E model is pre-trained on a 1B token subset of the SlimPajama-627B dataset (Soboleva et al., 2023). Moreover, the distributed training of the GPT-350M-16E model is experimented with the three different configurations, as shown in Table 2. Using the hybrid parallel strategy of ZeRO-2 DP (Rajbhandari et al., 2020) and EP, the training is deployed on a cluster that comprises a total of 60 nodes with 8×A800-SXM4-80GB GPUs each. The vision model, SwinV2-MoE (Hwang et al., 2023; Liu et al., 2021), is trained on the ImageNet-1K dataset (Deng et al., 2009).
@@ 6.2.1. Checkpoint Size
We evaluate the effectiveness of PEC in reducing checkpoint size through experiments on the GPT-350M-16E model training. Unless otherwise specified in subsequent experiments, PEC employs the sequential selection strategy, and the baseline refers to the method provided by the Megatron-DeepSpeed framework (Microsoft, 2022). As illustrated in Figure LABEL:fig:result_size_a, the total checkpoint size for each process decreases as $K_{pec}$ decreases, reaching 42.3% of the full model checkpoint size when $K_{pec}$ is set to 1.
However, merely reducing the total checkpoint size is insufficient for optimizing efficiency in distributed training. As the checkpointing workload is distributed across various training ranks, the duration of the blocking checkpointing process is primarily determined by the bottleneck rank, which has the heaviest workload and longest processing time. Therefore, we further assess the checkpointing workload of the bottleneck rank using different sharding strategies, as illustrated in Figure LABEL:fig:result_size_b-LABEL:fig:result_size_d.
The results indicate that our fully sharded checkpointing strategy, which applies equal sharding to both non-expert and expert parts, significantly reduces the workload of bottleneck rank in both full saving (12% to 28%) and PEC scenarios (22% to 29%). Notably, equal sharding of the expert part is only effective in scenarios with multiple EP groups (Case 3). With $K_{pec}=1$, adaptive sharding of the non-expert part can further reduce the workload by 3.7% to 6.1%.
@@ 6.2.3. Asynchronous Checkpointing
Given that our approaches have been validated to reduce the duration of the checkpointing process, we further assess the end-to-end optimization efficacy of our MoC-System, which implements an asynchronous checkpointing process. As illustrated in Figure 21, the fully optimized asynchronous process in our MoC-System (“MoC-Async”) can decrease the overhead of each checkpointing process ($O_{save}$) by 98.2% to 98.9% and accelerate each training iteration by 3.25 to 5.12 times in the three experimental cases, compared to the baseline using blocking checkpointing. $O_{save}$ refers to the additional time that surpasses the normal training processes, including “F&B” and “Update,” as indicated by the duration beyond the red dotted line in Figure 21.
@@
When our asynchronous checkpointing is applied without optimization by PEC and fully sharded techniques (“Base-Async”), it can overlap 97.9% of the checkpointing time in Case 2, as the snapshot duration is sufficiently short to achieve complete overlap. However, this method can only overlap 86.3% and 92.1% in Case 1 and Case 3, respectively, because their durations are too long to be fully overlapped. Employing all optimizations, “MoC-Async” can achieve 1.4% to 33.2% improvements over “Base-Async” in the three cases.
@@
In addition to reducing $O_{save}$, our “MoC-Async” achieves half the $I_{ckpt}$ compared to the “Base-Async” method, as it takes only half the time to complete the snapshot and persist process. For instance, it reduces $I_{ckpt}$ from 2.3 to 1.2 in Case 2. Consequently, our “MoC-Async” can minimize the overall checkpoint overhead $O_{ckpt}$.
Figure 21. Duration of a training iteration across three configurations, utilizing three checkpointing methods: (1) baseline, (2) “Base-Async,” which uses basic asynchronous checkpointing without our PEC and fully sharded checkpointing techniques, and (3) “MoC-Async,” representing the fully optimized asynchronous process within our MoC-System. “MoC-Async” can reduce checkpointing overhead by more than 98% compared to the baseline.
Figure 27. Scaling and generalizing results across various training factors. (a) Scaling the number of A800 GPUs using DP+EP parallelism, with each GPU assigned a unique expert for each MoE layer. (b) Scaling the number of A800 GPUs using DP+EP+TP parallelism, incorporating a 4-degree TP into the existing DP+EP configuration. (c) Scaling the number of H100 GPUs using DP+EP parallelism. (d) Generalizing different training sequence lengths. (e) Generalizing different model sizes. (f) The file size of the persist process across the DP+EP configurations with varying numbers of A800 GPUs.
6.2.4. Scaling and Generalizing
To demonstrate the scalability and generalizability of our proposed MoC-System, we conduct simulations to assess its efficacy across various configurations of training factors, including the number of GPUs, parallelism, hardware capability, sequence length, and model size. Our simulations utilize the ASTRA-SIM simulator (Rashidi et al., 2020) to model the computation and communication time costs for each iteration of distributed MoE model training across various scenarios with differing computing power and communication bandwidth. Aligned with our actual measured performance, we configure the A800 GPU simulations with 312 TFLOPS at a 20% utilization rate and a GPU-to-CPU snapshot bandwidth of 1 GB/s. Similarly, the H100 GPU simulations are set with 989 TFLOPS at a 20% utilization rate and a GPU-to-CPU snapshot bandwidth of 2 GB/s.
To reflect current large-scale training practices, we simulate the training of large LLaMA-like MoE models, which are commonly used in real-world applications (Jiang et al., 2024b; Liu et al., 2024a). In our simulations, depicted in Figure 27, we configure the MoE models with a hidden size of 2048, 16 attention heads, a head dimension of 128, an expert intermediate size of four times the hidden size, and 24 layers. Consistent with the three configurations used in Figure 21, Figure 27 illustrates the duration of a training iteration of three checkpointing methods: “Baseline,” “Base-Async,” and “MoC-Async.” Furthermore, Figure 27 breaks down the timing of the two asynchronous checkpointing methods to demonstrate the overlap duration of “F&B” and “Snapshot”, termed “F&B/Snapshot Overlap.”
Number of GPUs. To demonstrate the scalability of the MoC-System for large-scale training, we scale the model training across varying numbers of GPUs. Specifically, we employ Data Parallelism (ZeRO-2) and Expert Parallelism by assigning each expert of an MoE layer to a distinct GPU, scaling both the system and the model size. Figure LABEL:fig:scaling_a shows that the “F&B,” which can be used to overlap the snapshot overhead, significantly increases as the number of GPUs increases. Compared to the baseline, the two asynchronous checkpointing methods effectively facilitate overlap, thereby reducing time costs. When the number of GPUs is less than 1024, the snapshot duration of “Base-Async” is too long to be fully overlapped, resulting in lower efficiency compared to “MoC-Async.” In contrast, “MoC-Async,” configured to save only 1/8 of the experts per checkpoint, reduces the required snapshot time, making “F&B” the bottleneck of the total time when exceeding 64 GPUs. Additionally, “MoC-Async” can achieve substantial optimization even when “F&B” is considerably less than “Snapshot,” a benefit that “Base-Async” is unable to provide.
In this setup, an equal number of distinct expert parameters is allocated to each GPU, resulting in the data volume for each GPU-to-CPU snapshot remaining similar. However, as the number of GPUs increases, the data volume for CPU-to-storage persist on the cluster filesystem grows significantly, as illustrated in Figure LABEL:fig:scaling_f. Our “MoC-Persist” method significantly reduces the persist file size and the time required for the persist process, enabling shorter checkpoint intervals and consequently minimizing the lost time due to faults.
Parallelism. To generalize the efficacy of the MoC-System across various parallelism configurations, we further investigate training using the DP+EP+TP parallelism to train models with the same number of experts as in the DP+EP scenario, as depicted in Figure LABEL:fig:scaling_b. Although different parallelism strategies impact the iteration time of “F&B,” the behavior observed during checkpointing is similar to that seen with the DP+EP configuration (Figure LABEL:fig:scaling_a). Consistently, “MoC-Async” maintains optimal efficiency across all tested GPU configurations, particularly when the number of GPUs is fewer than 1024, where the snapshot duration of “Base-Async” cannot be fully overlapped.
Hardware Capability. To generalize the efficacy of the MoC-System across different hardware platforms, we conduct training simulations with the A800 GPU (Figure LABEL:fig:scaling_a) and the H100 GPU (Figure LABEL:fig:scaling_c) configurations. Due to variations in capabilities such as GPU computing power, GPU memory bandwidth, NVLink bandwidth, and GPU-to-CPU PCIe bandwidth, the durations of both “F&B” and “Snapshot” differ in the H100 scenario, resulting in varying overlap performance. Specifically, “MoC-Async” demonstrates significantly greater efficiency than other methods across all tested H100 configurations, as the snapshot of “Base-Async” cannot be fully overlapped even with 1024 GPUs. It is anticipated that the computation and communication capabilities associated with “F&B” will advance more rapidly than the GPU-to-CPU data transmission capabilities in future hardware platforms. Therefore, reducing the data volume of checkpointing through “MoC-Async” will remain valuable for larger-scale distributed training in the future.
Sequence Length. Sequence length is a critical factor influencing the “F&B” duration. To investigate its impact, we conduct simulations with varying sequence lengths using the DP+EP configuration on 256 A800 GPUs. While longer sequences significantly increase the “F&B” time, variations in sequence length do not impact the checkpointing process, as shown in Figure LABEL:fig:scaling_d. This is because the checkpointed data pertains to the constant model parameters rather than the dynamic activations. Consequently, “MoC-Async” can achieve higher efficiency across all evaluated sequence lengths.
Model Size. As larger model sizes lead to increased iteration times and larger data volumes for checkpointing, we conduct simulations using models of three different sizes: a hidden size of 1024 (“Small”), 2048 (“Medium”), and 3072 (“Large”), as illustrated in Figure LABEL:fig:scaling_e. These simulations are carried out using the DP+EP parallelism with 256 A800 GPUs. The results show that “MoC-Async” not only improves efficiency across various model sizes but provides more pronounced efficacy in scenarios involving larger-scale models. This is because model size impacts both the “F&B” and “Snapshot.” Due to the disparity in capabilities between computation and GPU-to-CPU data transmission, the duration of snapshots increases more significantly with the growth of the model size.